the length of a rectangle is 5 1/3 cm more than the width.the perimeter is 56 cm.Find the width of the rectangle.
Answer by macston(5194)

.
W=width; L=length=W+5 1/3
.
P=2(L+W)
56cm=2(L+W)
28cm=L+W
28cm=(W+5 1/3cm)+W
22 2/3cm=2W
11 1/3 cm=W
.
ANSWER: The width of the rectangle is 11 1/3 cm..
CHECK:
L=W+5 1/3 cm
L=11 1/3 cm+ 5 1/3 cm=16 2/3 cm
P=2(L+W)
56 cm = 2(16 2/3 cm+ 11 1/3 cm)
56 cm = 2(28cm)
56 cm = 56cm